# Global Deep Forecasting with Patient-Specific Pharmacokinetics

### v8

@@ A.1 Data
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
@@
Fig. 10 shows the first 10 days of data for 6 subjects from the simulated dataset. Fig. 11 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset. Figs. 12 and 13 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively. Table 4 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
Table 4: Sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set
@@ A.2 Models
The baseline models used in this work have been extensively used in previous blood glucose forecasting work. For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020) have been used in proposed models and as baselines.
@@
Exponential Smoothing (AutoETS) The AutoETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
@@
DLinear (DLinear) A Linear model that leverages linear layers to model the trend and seasonal components. A moving average filter is used to separate the time series into its trend and seasonal components (Zeng et al., 2023).
@@
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
@@
Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
@@
Neural Basis Expansion Analysis with Exogenous (NBEATSx) An MLP-based deep learning model that decomposes the input signal into trend and seasonality components using polynomial and harmonic basis projections. The NBEATSx extension enables modeling of exogenous variables in addition to the univariate signal (Oreshkin et al., 2020; Olivares et al., 2022a).
@@
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts (Elman, 1990; Cho et al., 2014).
@@
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
@@
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
@@
Informer (Informer) (Zhou et al., 2021) A Transformer-based deep learning model with a multi-head attention mechanism that uses autoregressive features from a convolution network. A ProbSparse self-attention mechanism and a self-attention distilling process are leveraged to reduce computational complexity.
@@
Patch time series Transformer (PatchTST) (Nie et al., 2023) A Transformer-based deep learning model with a multi-head attention mechanism. Input time series into sub-series level patches as input tokens.
@@
Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies. TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
@@ A.3 Model Training and Hyperparameters
The deep learning models were trained using an NVIDIA A100 Tensor Core GPU. For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 5. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 5. We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
@@
We tune additional specific hyperparameters for each architecture. For the MLP model, we tune the number of hidden units per layer ([128, 256]). For the LSTM model, we tune the hidden size ([128, 256]). For the TCN model, we tune the hidden size ([128, 256]) and number of dilations ([2, 4, 8]). For the PatchTST model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and patch length ([8, 16, 32, 64, 96, 128]). For the Informer model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and number of decoder layers ([1, 2]). For the TFT model, we tune the hidden size ([128, 256, 512]) and number of attention heads ([4, 8, 16]). For the NBEATSx and NHITS models, we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
A.4 Propositions
@@
[…]
@@
A.5 Algorithm
We outline the algorithm for training our proposed PK encoder that transforms sparse exogenous variables into time-dependent concentration profiles. This framework can be implemented in any deep learning architecture. It can also be adapted to utilize other concentration-time profile functions pertinent to a specific healthcare context.
The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch. This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library (Liaw et al., 2018). For each training step, the algorithm iterates over batches of data. It computes a context vector c based on the input data x, k, and other parameters. It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values. After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step. It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$. It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss. Finally, the algorithm returns the final k values.
Algorithm 1 PK Encoder Training.
$E$ is the number of trainings steps, $\eta$ is the learning rate, $N$ is the number of subjects, $H$ is the forecast horizon, and $L$ is the lag time.
Data: $N>0,f\in\mathbb{R}$, $L\in\mathbb{R}$, $H\in\mathbb{R}$
Result: k
$\mathcal{B}_{train}\leftarrow$ (split training data into batches)
$\mathcal{B}_{val}\leftarrow$ (split validation data into batches)
$\textbf{k}^{*}\leftarrow$ (Randomly select $\textbf{k}^{*}$ from defined hyperparmeter options, $\textbf{k}^{*}\in\{1,\ldots,2\}$)
@@ A.5 Algorithm
       $\textbf{k}\leftarrow\textbf{k}^{*}$triangleright Initialize k
for batch $b\in\mathcal{B}_{train}$ do
               $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
$\textbf{k}\leftarrow\textbf{k}-\eta\cdot\nabla_{\textbf{k}}\ell(\textbf{k})$triangleright Update k using SGD
@@
      if i % validation step == 0 then
               $\ell(\textbf{k})\leftarrow 0$
for batch $b\in\mathcal{B}_{val}$ do
                     $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
@@
            $\textbf{k}^{*}\leftarrow\arg\min_{\textbf{k}^{*}\in\{1,\ldots,2\}}\ell(\textbf{k})_{total}$triangleright Select the hyperparameter that minimizes loss
@@
return k
Figure 10: Time series for the first 10 days of 6 subjects from the simulated dataset.
Figure 11: Time series for the first 10 days of 6 subjects from the OhioT1DM dataset.
Figure 12: Density plots of bolus and basal insulin for 12 simulated data subjects. Bolus doses are general larger than basal doses.
Figure 13: Density plots of bolus and basal insulin for OhioT1DM subjects. Bolus doses are general larger than basal doses.
@@ C.1 Patient-specific (Local) Models
We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects. Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table. 6 and Fig. 6. Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig. 5.
Table 6: MAE computed across forecast horizon values using local and global models for the OhioT1DM dataset. Best results are in bold.
@@ C.2 Forecasting with Dose-Independent Parameters
Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180). Result for models with (w/) and without (w/o) exogenous variables are shown. Best results are in bold.
@@
w/o Exog. w/ Exog. w/o Exog. w/ Exog.
@@
We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values. However, this approach led to worse performance on the OhioT1DM dataset. The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters (Xie, 2018; Visentin et al., 2016). In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3. Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data. These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
C.3 Model MAE and RMSE Results
This appendix presents the complete results for all models for both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis. Tables 8 and 9 provide complete MAE and RMSE results, respectively.

### v9

@@ A.1 Data
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) patients with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
@@
Table 4: Sample count (% missing samples) for the original OhioT1DM datasets and preprocessed test set
@@ A.2 Models
Exponential Smoothing (AutoETS) - The AutoETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past (Hyndman, 2008).
@@
DLinear (DLinear) - A linear model that leverages linear layers to model the trend and seasonal components. A moving average filter is used to separate the time series into its trend and seasonal components (Zeng et al., 2023).
@@
Multi Layer Perceptrons (MLP) - A neural network architecture composed of stacked fully connected neural networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
@@
Neural Hierarchical Interpolation for Time Series (NHITS) - A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
@@
Neural Basis Expansion Analysis with Exogenous (NBEATSx) - An MLP-based deep learning model that decomposes the input signal into trend and seasonality components using polynomial and harmonic basis projections. The NBEATSx extension enables modeling of exogenous variables in addition to the univariate signal (Oreshkin et al., 2020; Olivares et al., 2022a).
@@
Recurrent Neural Network (RNN) - A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts (Elman, 1990; Cho et al., 2014).
@@
Long Short-Term Memory Recurrent Neural Network (LSTM) - A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts (Sak et al., 2014).
@@
Temporal Convolution Network (TCN) - A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
@@
Informer (Informer) - A Transformer-based deep learning model with a multi-head attention mechanism that uses autoregressive features from a convolution network. A ProbSparse self-attention mechanism and a self-attention distilling process are leveraged to reduce computational complexity (Zhou et al., 2021).
@@
Patch time series Transformer (PatchTST) - A Transformer-based deep learning model with a multi-head attention mechanism. The input time series is segmented into fixed-length sub-series-level patches, which serve as input tokens (Nie et al., 2023).
@@
Temporal Fusion Transformer (TFT) - An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies. TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
@@ A.3 Model Training and Hyperparameters
The deep learning models were trained using an NVIDIA A100 Tensor Core GPU. For all models, we use the hyperparameter values shown in Table 5 and tune the learning rate and random seed, in addition to architecture-specific hyperparameters discussed below. Additionally, for all models with the PK encoder, we tune the initial k values as shown in the second section of table 5. We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
@@
In addition to tuning the learning rate and random seed for each model, we also tune architecture-specific hyperparameters. For the DLinear model, we tune the moving average window size ([5, 15, 25, 35]). For the MLP model, we tune the number of hidden units per layer ([128, 256, 512]). For the LSTM and RNN models, we tune the hidden size ([128, 256, 512]). For the TCN model, we tune the hidden size ([128, 256, 512]) and number of dilations ([2, 4, 8]). For the PatchTST model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and patch length ([8, 16, 32, 64, 96, 128]). For the Informer model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and number of decoder layers ([1, 2]). For the TFT model, we tune the hidden size ([128, 256, 512]) and number of attention heads ([4, 8, 16]). For the NBEATSx and NHITS models, we did not tune additional hyperparameters. The code to train models and tune hyperparameters for each architecture is included in the paper’s code repository.
A.4 Model Evaluation
Metrics.
We evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE),
Here, $\mathbf{y}_{\tau}$ are the blood glucose values, and $\hat{\mathbf{y}}_{\tau}$ are the model predictions for the forecast horizon $H$, with $\tau=\{T+1,\dots,T+H\}$. We compute the metrics across all points within the forecast horizon and present the average results over 8 trials of independently trained models.
Statistical Significance Tests.
To evaluate the performance differences between our hybrid global-local PK model and the baseline models, we conducted statistical significance testing on the forecasting MAE (Mean Absolute Error) results. For each statistical comparison, patient results were averaged over eight trials of individually trained models, and the paired t-test was computed by pairing patient results between the hybrid global-local PK model and the baseline model. The null hypothesis for each test was that there is no difference in performance between our hybrid global-local PK model and the compared baseline across patients. We considered results statistically significant at the conventional threshold of $\text{p-value}<0.05$, indicating strong evidence against the null hypothesis.
A.5 Propositions
@@
$x$ $\propto\int_{0}^{T}C(t,\cdot)_{x}dt$ (15)
$n\cdot x$ $=n\int_{0}^{T}C(t,\cdot)_{x}dt$ (16)
$=\int_{0}^{T}nC(t,\cdot)_{x}dt$ (17)
$=\int_{0}^{T}\underbrace{C(t,\cdot)_{x}+C(t,\cdot)_{x}+\cdots+C(t,\cdot)_{x}}_{n\text{ times}}dt$ (18)
@@
A.6 Algorithm
We present the training procedure for the hybrid global-local PK model in Algorithm 1. This framework is compatible with any deep learning architecture capable of modeling exogenous variables alongside the endogenous target forecast signal.
The algorithm initializes k as a vector of network embedding weights with values between 1 and 2 before model training. The initial value of k is treated as a hyperparameter and tuned using the Hyperopt package from the RayTune library (Liaw et al., 2018). At each training step, the algorithm iterates over batches of data. The model $f_{\theta}$ generates forecasts $\hat{\textbf{y}}_{\tau}$ based on the input data. Network parameters, including k, are updated using stochastic gradient descent (SGD) based on the loss of the model $\ell(f_{\theta})$, computed between the predicted values $\hat{\textbf{y}}$ and the true values y using the loss function $\mathcal{L}$. After a pre-specified number of training steps, the algorithm performs a validation step by calculating the total validation loss $\ell(f_{\theta})$. The final hyperparameters are selected based on the configuration that minimizes validation loss across all runs and are used to train the final model. The algorithm returns the final trained model $f_{\theta}$.
Algorithm 1 Hybrid Global-Local PK Model Training.
$E$ is the number of training steps, $\eta$ is the learning rate, $N$ is the number of subjects, $\nu$ is the sampling interval of the data, $L$ is the lag time, and $H$ is the forecast horizon. The model $f_{\theta}$ hyperparameters include the learning rate $\eta$, the hidden size h, and the initial (init) k values for bolus, basal, and CHO concentration curves, among other hyperparameters. For brevity, we define $\textbf{k}_{\text{init}}=(\textbf{k}_{\text{init, bolus}},\textbf{k}_{\text{init, basal}},\textbf{k}_{\text{init, CHO}})$ to represent the initial concentration curve parameters. $\mathcal{L}$ is the loss function (Huber Loss). $V$ is the specified validation step, or the number of training steps between each validation loss check. Forecasts $\hat{\textbf{y}}$ are generated by $f_{\theta}$ as a function of historical glucose values y, treatment concentration curves $\widecheck{\textbf{x}}$, and static features s.
Input: $N>0$, $\nu>0$, $L>0$, $H>0$, $\eta>0$, $\textbf{h}\in\{128,256,512,1024\}$, $\textbf{k}_{\text{init}}\in[1,2]^{N\times 1}$
Output: $f_{\theta}$
$\mathcal{B}_{\text{train}}\leftarrow$ (split training data into batches)
$\mathcal{B}_{\text{val}}\leftarrow$ (split validation data into batches)
for each run $o$ in hyperparameter tuning samples (runs) $O$ do
    $f_{\theta}\leftarrow$ (initialize the model)
$\textbf{k}\leftarrow\textbf{k}_{\text{init}}$ $\triangleright$ Each entry in k is initialized to the same scalar value
$\ell(f_{\theta})_{\text{val}}^{(o)}\leftarrow 0$ $\triangleright$ Initialize validation loss for run $o$
@@ A.6 Algorithm
    for batch $b\in\mathcal{B}_{\text{train}}$ do
          $\widecheck{\textbf{x}}\leftarrow\sum_{t=0}^{L}C(\nu\textbf{w}_{t,:},\textbf{x}_{t},\textbf{k})$
$\hat{\textbf{y}}_{\tau}\leftarrow f_{\theta}(\textbf{y},\widecheck{\textbf{x}},\textbf{s})$
$\ell_{\text{batch}}(f_{\theta})\leftarrow\mathcal{L}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $\triangleright$ $\tau$ represents the forecast horizon
$\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\ell_{\text{batch}}(f_{\theta})$ $\triangleright$ update network parameters using SGD
@@
    if $i\mod V==0$ then
          $\ell(f_{\theta})_{\text{total}}\leftarrow 0$
for batch $b\in\mathcal{B}_{\text{val}}$ do
             $\widecheck{\textbf{x}}\leftarrow\sum_{t=0}^{L}C(\nu\textbf{w}_{t,:},\textbf{x}_{t},\textbf{k})$
$\hat{\textbf{y}}_{\tau}\leftarrow f_{\theta}(\textbf{y},\widecheck{\textbf{x}},\textbf{s})$
$\ell_{\text{batch}}(f_{\theta})\leftarrow\mathcal{L}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$
$\ell(f_{\theta})_{\text{total}}\leftarrow\ell(f_{\theta})_{\text{total}}+\ell_{\text{batch}}(f_{\theta})$
@@
       $\ell(f_{\theta})_{\text{val}}^{(o)}\leftarrow\frac{\ell(f_{\theta})_{\text{total}}}{|\mathcal{B}_{\text{val}}|}$ $\triangleright$ Average loss over entire validation set and update loss
@@
end for
$\{\eta,\textbf{h},\textbf{k}_{\text{init}},\ldots\}\leftarrow\arg\min_{o\in O}\ell(f_{\theta})_{\text{val}}$ $\triangleright$ select best hyperparameters; retrain final model
return $f_{\theta}$
Figure 11: Time series for the first 10 days of 6 patients from the simulated dataset.
Figure 12: Time series for the first 10 days of 6 patients from the OhioT1DM dataset.
Figure 13: Density plots of bolus and basal insulin for 12 simulated data patients. Bolus doses are general larger than basal doses.
Figure 14: Density plots of bolus and basal insulin for OhioT1DM patients. Bolus doses are general larger than basal doses.
@@ C.1 Patient-specific (Local) Models
We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects. Global PK models consistently outperform local PK models, on average across 8 trials, for each OhioT1DM patient as shown in Table 6 and Fig. 6. Furthermore, the global NHITS-PK model outperforms the global NHITSsparse exogenous model for 11 out of 12 OhioT1DM patients, on average across 8 trials, as shown in Table 6.
Table 6: MAE computed across forecast horizon values using local and global models for the OhioT1DM dataset. The best results are in bold.
@@ C.2 Forecasting with Dose-Independent Parameters
Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose $\leq 70|\geq 180$). Result for models with (w/) and without (w/o) exogenous variables are shown. The best results are in bold.
@@
w/o Exog.
w/ Exog.
w/o Exog.
w/ Exog.
@@
We found that learning a single k value for both bolus and basal insulin improved performance on the simulated dataset only when forecasts were evaluated at critical values. However, this approach resulted in worse performance when evaluated across all values in the simulated dataset, as well as on the OhioT1DM dataset. The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static parameters for insulin-glucose kinetics (Xie, 2018; Visentin et al., 2016). In contrast, insulin may exhibit nonlinear dose-dependent PK in real-world scenarios, as mentioned in section 3.3.